OpenStudy (anonymous):

Find the value of X... Also, Could someone tell me when you use sin-1 and sin( ?

OpenStudy (anonymous):



OpenStudy (anonymous):

sin = opposite / hyptoenuse sin 38 = 17/x Switch them: 17/x = sin 38 Use a calculator to find sin 38 17/x = .296 Now, mulitply by 17 to get x alone x = 5.04 Hope this helps :o)

OpenStudy (anonymous):

When I do sin38 I get .615

OpenStudy (anonymous):

You're right

OpenStudy (anonymous):

Oops, you're right. Lemme redo that part. I had my calculator in a different mood, ops! :o)

OpenStudy (anonymous):

17/x = .615 x = 10.47